OpenStudy (anonymous):

solve for x 6.2(x-15.1)=9.3(-x-3.2)

OpenStudy (anonymous):

first you need to expand the brackets 6.2(x-15.1)=6.2x-93.62 9.3(-x-3.2)=-9.3x-29.76

OpenStudy (anonymous):

6.2x-93.62=-9.3x-29.76 Than take the x parts to one side and the number bits to the other So add 9.3x to both sides 6.2x+9.3x-93.62=-9.3x+9.3x-29.76

OpenStudy (anonymous):

That is 15.5x-93.62=0-29.76 Now add 93.62 15.5x=93.62-27.76=65.86

OpenStudy (anonymous):

Start by distributing... 6.2x - 96.62 = -9.3x - 29.76 Then you get x by itself 15.5x = 66.86 x = 4.31

OpenStudy (anonymous):

Now divide by 15.5 to get the result x=65.86/15.5=4.249

OpenStudy (anonymous):

If you are unsure that you have the right result you can check it by putting it back to the original equation: 6.2(x-15.1)=9.3(-x-3.2) 6.2(4.249-15.1)=9.3(-4.249-3.2) -67.2762=-69.2757 That is not true so the result is not right :-) 6.2(4.31-15.1)=9.3(-4.31-3.2) -66.898=-69.843 :DDDD

OpenStudy (anonymous):

Ok i give up :-) I hope you get the idea

OpenStudy (anonymous):

Andras, Thank you but if it doesn't check isn't wrong?

OpenStudy (anonymous):

Somewhere I guess I miscalculated something. Happens often after working 10 hours

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

GsSpade23, Thank you. I am not sure if that is correct.....